# Simulating Polynomial-Time Nondeterministic Turing Machines via Nondeterministic Turing Machines

### v21

@@ 7. Rich Structure of ${\rm co}\mathcal{NP}$
It is sufficient to show that for every language $L\in{\rm co}\mathcal{NP}$, $L\leq_{p}{\rm TAUT}$, where $\leq_{p}$ is polynomial-time many-one reduction [Kar72]. We can modify the Cook-Levin reduction [Coo71, Lev73] from $\overline{L}=\{0,1\}^{*}-L$ to $\rm SAT$} (recall that all nondeterministic Turing machines are with input alphabet $\{0,1\}$ because that will be all we need; see Section 3). So for any input $w\in\{0,1\}^{*}$, the Cook-Levin reduction produces a formula $\varphi_{w}$ that is satisfiable if and only if $w\in\overline{L}$. In other words, the formula $\neg\varphi_{x}$ is in $\rm TAUT$} if and only if $x\in L$, which completes the proof.
Now, we are naturally at a right point to give the proof of Theorem 1.6 as follows:
Proof of Theorem 1.6. Let $L=\{0,1\}^{*}-L_{\rm inter}$, where the language $L_{\rm inter}$ is the same as in Lemma 7.1. Such a language indeed exists by Corollary 1.2, which states that $\mathcal{P}\neq\mathcal{NP}$.
@@ 8. Frege Systems
The focal point of this section is Frege systems, which are very strong proof systems in a propositional setting [CR79], based on axiom schemes and rules such as modus ponens. While Frege systems operate with Boolean formulas as lines, the extended Frege system EF works with Boolean circuits; see e.g., [Jer05]. Furthermore, showing lower bounds on Frege or even extended Frege systems constitutes a major open problem in proof complexity; see e.g., [BP01].
@@
In particular, Frege proof systems are proof systems for propositional logic. As a matter of fact, Frege systems are the usual‘ ‘textbook” proof systems for propositional logic based on axioms and rules; see e.g., [CR79]. A Frege system composed of a finite set of axiom schemes and a finite number of rules is a possible axiom scheme. Furthermore, a Frege proof is a sequence of formulas where each formula is either a substitution instance of an axiom or can be inferred from previous formulas by a valid inference rule. At the same time, Frege systems are required to be sound and complete. However, the exact choice of the axiom schemes and rules does not matter, as any Frege systems are polynomially equivalent; see e.g., [CR79] or see Theorem 8.1 below. Thus, we can assume without loss of generality that modus ponens (see e.g., Remark 8.3 below) is the only rule of inference.
@@ A Frege proof system $F$ is an inference system for propositional logic based on
The axiom schemes of a Frege proof system is the following (see e.g., [Bus99]):
@@
An abstract propositional proof system is a polynomial-time computable function $g$ such that $Range(g)=TAUT$; i.e., the range of $g$ is the set of all Boolean tautologies. An $g$-proof of a formula $\phi$ is a string $w$ such that $g(w)=\phi$.
@@ 8.1. Lower Bounds for Frege Systems
The following theorem gives a necessary and sufficient condition for the complexity class $\mathcal{NP}$ being closed under complementation. For completeness, the proof is also quoted from [CR79]:
@@ 9. Open Problems
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contain the problem Satisfiability? Note that by footnote 1, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this can not exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. If one could show that the problem of Satisfiability is not in $\mathcal{BQP}$, then the relationship between $\mathcal{BQP}$ and $\mathcal{NP}$ would be clear. Currently, we have no additional information about this problem. We tried to think about it but also failed, and we remark that this is also a fascinating open problem in complexity theory.
@@ Appendix B A New Construction of $L_{s}^{{}^{\prime}i}$
    $U$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}$ machine $M_{l}$ for some $l$ (by Definition 2.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic Turing machine, i.e., $x$ represents the trivial Turing machine with an empty next-move function), then GOTO $5$, else determine $t$, the number of tape symbols used by $M_{l}$; $s$, its number of states; and $k$, its order. The third tape of $U$ can be used as “scratch” memory to calculate $t$.
@@
    If $k+1\leq i+1$, then GOTO $3$, else reject the input $x$.
@@
Compared with the proof of Theorem 4.1, the construction of $U$ is only different in the red item 2.

### v22

@@ 7. Rich Structure of ${\rm co}\mathcal{NP}$
It is sufficient to show that for every language $L\in{\rm co}\mathcal{NP}$, $L\leq_{p}{\rm TAUT}$, where $\leq_{p}$ is polynomial-time many-one reduction [Kar72]. We can modify the Cook-Levin reduction [Coo71, Lev73] from $\overline{L}=\{0,1\}^{*}\setminus L$ to $\rm SAT$} (recall that all nondeterministic Turing machines are with input alphabet $\{0,1\}$ because that will be all we need; see Section 3). So for any input $w\in\{0,1\}^{*}$, the Cook-Levin reduction produces a formula $\varphi_{w}$ that is satisfiable if and only if $w\in\overline{L}$. In other words, the formula $\neg\varphi_{x}$ is in $\rm TAUT$} if and only if $x\in L$, which completes the proof.
Now, we are naturally at the right point to give the proof of Theorem 1.7 as follows:
Proof of Theorem 1.7. Let $L=\{0,1\}^{*}\setminus L_{\rm inter}$, where the language $L_{\rm inter}$ is the same as in Lemma 7.1. Such a language indeed exists by Corollary 1.2, which states that $\mathcal{P}\neq\mathcal{NP}$.
@@ 8. Frege Systems
The focal point of this section is Frege systems, which are very strong proof systems in a propositional setting [CR79], based on axiom schemes and rules such as modus ponens. While Frege systems operate with Boolean formulas as lines, the extended Frege system EF works with Boolean circuits; see, e.g., [Jer05]. Furthermore, showing lower bounds on Frege or even extended Frege systems constitutes a major open problem in proof complexity; see e.g., [BP01].
@@
In particular, Frege proof systems are proof systems for propositional logic. As a matter of fact, Frege systems are the usual “textbook” proof systems for propositional logic based on axioms and rules; see, e.g., [CR79]. A Frege system composed of a finite set of axiom schemes and a finite number of rules is a possible axiom scheme. Furthermore, a Frege proof is a sequence of formulas where each formula is either a substitution instance of an axiom or can be inferred from previous formulas by a valid inference rule. At the same time, Frege systems are required to be sound and complete. However, the exact choice of the axiom schemes and rules does not matter, as any Frege systems are polynomially equivalent; see e.g., [CR79] or see Theorem 8.1 below. Thus, we can assume without loss of generality that modus ponens (see e.g., Remark 8.3 below) is the only rule of inference.
@@ A Frege proof system $F$ is an inference system for propositional logic based on
The axiom schemes of a Frege proof system are the following (see e.g., [Bus99]):
@@
An abstract propositional proof system is a polynomial-time computable function $g$ such that $Range(g)=TAUT$; i.e., the range of $g$ is the set of all Boolean tautologies. A $g$-proof of a formula $\phi$ is a string $w$ such that $g(w)=\phi$.
@@ 8.1. Lower Bounds for Frege Systems
The following theorem gives a necessary and sufficient condition for the complexity class $\mathcal{NP}$ being closed under complement. For completeness, the proof is also quoted from [CR79]:
@@ 9. Open Problems
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contain the problem Satisfiability? Note that by Corollary 1.4, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this cannot exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. So, if one is able to show that the problem Satisfiability is not in $\mathcal{BQP}$, then the precise relationships between $\mathcal{BQP}$ and $\mathcal{NP}$ would be clear. Currently, we have no additional information about this problem. We tried to think about it but also failed, and we remark that this is also a fascinating open problem in complexity theory.
@@ Appendix B A New Construction of $L_{s}^{{}^{\prime}i}$
    $U$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}$ machine $M_{l}$ for some $l$ (by Definition 2.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic Turing machine, i.e., $x$ represents the trivial Turing machine with an empty next-move function), then GOTO $5$; else determine $t$, the number of tape symbols used by $M_{l}$; $s$, its number of states; and $k$, its order. The third tape of $U$ can be used as “scratch” memory to calculate $t$.
@@
    If $k+1\leq i+1$, then GOTO $3$; else reject the input $x$.
@@
Compared with the proof of Theorem 4.1, the construction of $U$ is only different in the item 2.